OpenStudy (anonymous):

someone please help me with some algebra 1 problems??? medal???

OpenStudy (anonymous):

1. \[4\sqrt{6} \times 6\sqrt{2}\]

OpenStudy (anonymous):

@UnkleRhaukus

OpenStudy (anonymous):

@ranga

OpenStudy (anonymous):

someone?!

OpenStudy (unklerhaukus):

what are the factors of 6?

OpenStudy (anonymous):

i figured it out! thank you(=